# Positivity bounds from thermal field theory entropy

### v1

@@ Abstract
We present a novel approach to deriving positivity bounds on effective field theories by analyzing the thermodynamic behavior of thermal quantum field systems. Focusing on scalar theories with higher-dimensional operators, we compute the finite-temperature entropy using thermal field theory techniques. We argue that consistency with fundamental thermodynamic principles—specifically, the expectation that entropy increases with the introduction of new degrees of freedom—imposes nontrivial constraints on Wilson coefficients. In particular, we show that the coefficient of the leading dimension-8 operator must be strictly positive. This thermodynamic perspective offers an alternative to traditional S-matrix-based derivations of positivity bounds and provides fresh insight into the interplay between entropy, unitarity, and causality in quantum field theory.
@@ I Introduction
In this work, we explore an alternative approach. Motivated by previous studies connecting the thermodynamic properties of black hole entropy to quantum gravity conjectures [57, 58, 59, 60, 61, 62, 63, 64, 65, 66], we investigate whether similar principles can be applied within the framework of finite-temperature quantum field theory. Our central proposal is that enforcing thermodynamic consistency—specifically, requiring that the thermal entropy does not decrease when additional degrees of freedom are introduced—naturally leads to positivity constraints on the Wilson coefficients in the EFT expansion.
To make this idea concrete, we analyze a scalar field theory with a shift symmetry, compute its entropy at finite temperature, and show that ensuring a physically meaningful entropy implies a strictly positive coefficient for the leading higher-dimensional operator. This suggests a deep connection between statistical mechanics and the consistency structure of effective field theories, offering a complementary perspective to traditional amplitude-based arguments.
@@ Acknowledgements
We are grateful to Anna Tokareva for carefully reading the manuscript and for insightful comments. Y.X. also thanks Longqi Shao and Shuang-Yong Zhou for inspiring discussions.

### v2

@@ Abstract
We present an approach to deriving positivity bounds on effective field theories by analyzing the thermodynamic behavior of thermal quantum field systems. Focusing on scalar theories with higher-dimensional operators, we compute the finite-temperature entropy using thermal field theory techniques. We argue that consistency with fundamental thermodynamic principles—specifically, the expectation that entropy increases with the introduction of new degrees of freedom—imposes nontrivial constraints on Wilson coefficients. In particular, we show that the coefficient of the leading dimension-8 operator must be strictly positive. This thermodynamic perspective offers an alternative to traditional S-matrix-based derivations of positivity bounds and provides a complementary perspective into the interplay between entropy, unitarity, and causality in quantum field theory.
@@ I Introduction
In this work, we explore an alternative approach. Motivated by previous studies connecting the thermodynamic properties of black hole entropy to quantum gravity conjectures [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67], we investigate whether similar principles can be applied within the framework of finite-temperature quantum field theory. Our central proposal is that enforcing thermodynamic consistency—specifically, requiring that the thermal entropy does not decrease when additional degrees of freedom are introduced—naturally leads to positivity constraints on the Wilson coefficients in the EFT expansion.
To make this idea concrete, we analyze a scalar field theory with a shift symmetry, compute its entropy at finite temperature, and show that ensuring a physically meaningful entropy implies a strictly positive coefficient for the leading higher-dimensional operator. This suggests a deep connection between statistical mechanics and the consistency structure of effective field theories, offering a complementary route to traditional amplitude-based arguments.
@@ V Conclusion and discussion
After posting v1 of this manuscript, we became aware of a closely related preprint by Fernández-Sarmiento, Penco, and Rosen [72] through correspondence from Riccardo Penco. Their analysis formulates an entropy-positivity condition in terms of a negative thermal shift of the grand potential and tests it across several systems, together with a discussion of regimes where the bound may not apply. Our approach starts from a different argument: by beginning with the Araki-Lieb inequality and an explicit leading-order thermal QFT computation, we demonstrate this condition in a Lorentz-invariant scalar EFT and obtain the positive sign of the leading dimension-8 operator. We also outline the assumptions of our derivation and its regime of validity within this setup.
@@ Acknowledgements
We are grateful to Anna Tokareva for a careful reading of the manuscript and for insightful comments. Y.X. also thanks Longqi Shao and Shuang-Yong Zhou for inspiring discussions, and colleagues at the Amplitudes 2025 conference (Seoul National University, June 16–20, 2025) for helpful feedback on a preliminary version of this work presented as a poster. Finally, we thank Riccardo Penco for the correspondence and for bringing their nice paper to our attention.